OpenStudy (anonymous):

Find the derivative of the function using the definition of derivative. f(x) = (8 + x )/(1 − 8x)

OpenStudy (aum):

\[ f'(x) = \lim_{h \rightarrow 0}\frac{f(x+h) - f(x)}{h} ~~~~~~~ \text{----- (1)}\\ f(x) = \frac{8+x}{1-8x} \\ f(x+h) = ? \\ \text{ } \\ \text{Plug f(x+h) and f(x) into (1), simplify and take the limit} \]

OpenStudy (anonymous):

How do I simplify it? That is what I'm having difficulty with.

OpenStudy (aum):

\[ f'(x) = \lim_{h \rightarrow 0}\frac{f(x+h) - f(x)}{h} ~~~~~~~ \text{----- (1)}\\ f(x) = \frac{8+x}{1-8x} \\ f(x+h) = \frac{8+x+h}{1-8x-8h} \\ f(x+h) - f(x) = \frac{8+x+h}{1-8x-8h} - \frac{8+x}{1-8x} = \\ \frac{(8+x+h)(1-8x)-(8+x)(1-8x-8h)}{(1-8x-8h)(1-8x)} = \\ \frac{(8+x+h)(1-8x)-(8+x)(1-8x-8h)}{(1-8x-8h)(1-8x)} = \\ \frac{\cancel{8}-\cancel{64x}+\cancel{x}-\cancel{8x^2}+h-\cancel{8xh}-\cancel{8}+\cancel{64x}+64h-\cancel{x}+\cancel{8x^2}+\cancel{8xh}}{(1-8x-8h)(1-8x)} = \\ \frac{65h}{(1-8x-8h)(1-8x)} \\ f'(x) = \lim_{h \rightarrow 0}\frac{f(x+h) - f(x)}{h} = \lim_{h \rightarrow 0}\frac{65h}{h(1-8x-8h)(1-8x)} = \\ \lim_{h \rightarrow 0}\frac{65}{(1-8x-8h)(1-8x)} = \frac{65}{(1-8x)(1-8x)} = \frac{65}{(1-8x)^2} \]